OpenStudy (anonymous):

[6.02] For the following system, if you isolated x in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 3x + y = 8 -x - 2y = -10

OpenStudy (anonymous):

i really need help i have half hour to turn this is please help me

OpenStudy (anonymous):

If you isolated "X" in the second equation, you would take the answer you got (10) and plug it into the first equation to find "Y". which would give you y=4/15 Solution: x=10 from second equation y=4/15 from first equation i think thats what your looking for, if not im not sure how much help i'll be, sorry ):